OpenStudy (abbycross167):

Can someone please check the answer to my question? Please? If 3x + 4y = 9, then x equals a. 3-4y b. 9-4y c. 9+4y 3 d. 9-4y 3

OpenStudy (anonymous):

D

OpenStudy (anonymous):

dude dont blert out the answer

OpenStudy (imsleeziboii):

explain please :)

OpenStudy (anonymous):

3x-4y = 7 is your line We need its slope. Congruent slopes are PARALLEL. [Negative reciprocal slopes are PERPENDICULAR.] (1) Convert to y=mx+b (slope intercept form) to find its slope. 3x-4y = 7 -4y = 7-3x [move over the 3x] y = -(7/4) + (3/4)x [divide both sides by -4] y = (3/4)x - (7/4) [rearrange] WHAT IS THIS? IT IS A LINE WITH A SLOPE OF 3/4 AKA 0.75 AND A Y-INTERCEPT OF -(7/4) (2) Your given line has a slope of 3/4 aka 0.75. Your PARALLEL line will ALSO have a slope of 3/4 aka 0.75 (3) Plug your point (x,y) which was (-4,-2) and your known slope (3/4) into y=mx+b to solve for b (y-intercept) y = mx+b -2 = (3/4)(-4) + b [Substitute] -2 = -3 + b [multiply] b = 1 [Add 3 to right and left sides to isolate b] (4) REWRITE LINE WITH SLOPE (m) AND Y-INTERCEPT (b) m = 3/4 b = 1 y = mx + b y = (3/4)x + 1 y = 0.75x + 1 is also an acceptable answer

OpenStudy (anonymous):

there is your answer

OpenStudy (anonymous):

wow @silveralchemist09

OpenStudy (anonymous):

subtract 4 from both side then divide each side by 3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

It ask what is Y

Atsie (atsie):

@silveralchemist09 And THATS how its done! Good job! You explain very well!

OpenStudy (anonymous):

your welcome

OpenStudy (abbycross167):

Thank you all for helping me!!

OpenStudy (anonymous):

hey no problem

alones (alones):

@silveralchemist09 thanks for supporting OS rules <3

OpenStudy (anonymous):

youre welcome guys

OpenStudy (imsleeziboii):

silveralchemist just copied and pasted lol

alones (alones):

but.he didnt' give the direct answer out like someof you._.

OpenStudy (anonymous):

yah but i did the work i just saved it encase anyone else needs to use it

OpenStudy (imsleeziboii):

All i said was he copied and pasted, i told the guy to show the work too

OpenStudy (anonymous):

ok thanks

alones (alones):

Okay let's not exlode this..since we already got the answer with steps to it.

OpenStudy (anonymous):

yah youre right bye

OpenStudy (anonymous):

and thanks for medal